OpenStudy (anonymous):

Solve the following proportion for X? X/4 = 8/3 Round your answer to the nearest tenth. I forgot how to do this, help please?

OpenStudy (anonymous):

\[\frac{ x }{ 4 }=\frac{ 8 }{ 3 }\]

OpenStudy (anonymous):

3x=32 3x/3=x, 32/3=10.666666...

OpenStudy (anonymous):

I believe that's how you do it.

OpenStudy (anonymous):

(: Thank you. That helps

OpenStudy (anonymous):

No problem.